OpenStudy (anonymous):

what is the volume of a cube with a side length of 5cm?

OpenStudy (anonymous):

Volume of cube is equal to the cube of the side V=s^3 V=5^3 V=125 cm^3

OpenStudy (dmb1215):

yes that is true